# Smurfs: Leveraging Multiple Proficiency Agents with Context-Efficiency for Tool Planning

### v1

@@ 3.1 Framework Overview
The Executor Agent is responsible for choosing and executing the tools to solve the sub-tasks. The agent has access to an external tool library Qin et al. (2024). At each steps, the agent can invoke one tool to tackle the given task. As outlined in Algorithm 1, the agent, using the ReACT format Yao et al. (2022) to choose the tool and arguments, then execute the tool. More detailed information of the Executor Agent can be found in Appendix B.2.
Input: A task $q$, a hint from the verifier agent $h$, problem solving history $H$ and a set of available tools $T$.
Output: Tool response that is useful for solving the task.
thought $\leftarrow$ gen_thought($q$, $h$, $T$);
tool $\leftarrow$ choose_tool($q$, thought, $T$);
args $\leftarrow$ gen_arguments($q$, tool, $H$, tool_doc);
response $\leftarrow$ call_tool(tool, args);
Algorithm 1 Tool Call
@@
To mitigate the performance degradation caused by lengthy contexts, we introduce the ‘Answer Agent‘ role, designed to extract crucial content for each step and subtask. As demonstrated in the pilot study presented in Section 2.1, retaining all information may not always be beneficial, particularly in cases where the solution path is challenging to discern. Therefore, the primary role of the Answer Agent is to succinctly summarize the generated answers.
@@
Similar to current reasoning frameworks Wei et al. (2022); Yao et al. (2022), sequential task reasoning may lead to high computational and tool resource waste within the framework Qin et al. (2024). Therefore, inspired by Qin et al. (2024), we employ a depth-first search with an early stopping strategy to find the solution path. The Verifier Agent serves as an early-stopping and reflection mechanism in the intelligent multi-agent ensemble, allowing for a balance between effectiveness and efficiency. This mechanism not only ensures the accuracy of the generated responses but also optimizes resource use by preventing unnecessary computations. Moreover, if the Verifier Agent thinks the answer at this step isn’t accurate and reasonable enough, it will provide hints for the Executor Agent to get the missing information for the next steps. This dual role of the Verifier Agent enhances the overall performance of the framework, making it more robust and reliable for handling complex tasks.
3.2 Subtask Solving Process
After introducing the function of each agent, this section outlines how the agents collaborate to solve sub-tasks, as shown in Figure 3. Upon receiving a subtask or entering a new step, the tool list is refreshed, allowing the use of all available tools. The Executor Agent then calls the tools according to the task instruction. If a tool call fails, that tool is marked as unusable for this task at this step and the Executor Agent will try using other tools to solve the subtask at this step. This system also introduces a backtracking mechanism similar to DFSDT Qin et al. (2024) to handle the situation where errors frequently occur. Once the executor agent gets a correct response from a tool, the Answer Agent will refine the information, filtering out irrelevant details such as lengthy web pages and generate the answer for the subtask at this step using the local memory. After this, the Verifier Agent checks the answer’s accuracy. If it’s incorrect, the process returns to the Executor Agent with the hint from the Verifier Agent for the next step; Otherwise, the final answer is provided. If the Executor Agent reaches its retry limit without success, Answer Agent will review the entire process from the global memory to produce an answer. More details of the subtask solving process can be seen at  B
@@ 4 Experiments
To evaluate the effectiveness and efficiency of the Smurfs framework, we carried out a series of thorough experiments. In addition to the main experiment designed to assess the entire framework, we conducted an ablation study to test the capabilities of each component within the multi-agent framework. This section offers a detailed description of the experimental setup, methodologies employed, and key findings. Our goal is to showcase how the Smurfs framework, through the cooperative work of its multi-agent ensemble, effectively manages complex tasks while optimizing resource use. The experiments were designed with the following research objectives:
  • To validate the capability of the entire framework in managing tool planning tasks;
  • To independently assess the impact of each component on the overall performance, and identify the key factors influencing the multi-agent framework;
4.1 Evaluation
@@
Our experiments are conducted on ToolBench Qin et al. (2024), which encompasses multi-step tool usage tasks across over 16,000 APIs. To evaluate the planning and reasoning capabilities of the LLMs, we focused our experiments on intra-category multi-tool instructions (I2) and intra-collection multi-tool instructions (I3). These instructions involve selecting 2-5 tools from the same category or collection and sampling up to 3 APIs from each tool to formulate the instructions. We employed two metrics for evaluation: (1) Pass Rate measures the percentage of instructions successfully executed within the allocated budget, evaluated by ChatGPT. (2) Win Rate represents the preference selection by a ChatGPT evaluator when presented with two solution paths. All other settings are kept consistent with those of the ToolBench benchmark.
Models Method I2-Inst. I2-Cat. I3-Inst. Average
Pass Win Pass Win Pass Win Pass Win
ToolLLaMA-7B ReACT 30.5 50.8 31.5 41.8 25.0 55.0 29.0 49.2
DFSDT 77.0 68.5 77.0 58.0 66.0 69.0 73.3 65.2
Vicuna-7B ReACT & DFSDT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Smurfs (ours) 77 73 70.5 64.25 78.0 87.0 75.2 74.8
Mistral-Instruct-7B ReACT & DFSDT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Smurfs (ours) 77.5 80.0 79.5 79.2 79.0 94.0 78.7 84.4
GPT4 ReACT 67.0 65.8 72.0 60.3 47.0 78.0 62.0 68.0
DFSDT 79.5 73.3 77.5 63.3 71.0 84.0 76.0 73.5
Smurfs (ours) 71.0 77.5 72.0 77.0 64.0 89.5 69.0 81.3
Table 3: ToolBench evaluation, with some results derived from Qin et al. (2024); Yuan et al. (2024). The most effective approach is highlighted in bold, while the second is underlined.
4.2 Baselines
To investigate the varying impacts of the agent framework on models with different capabilities, we categorize our baseline into three groups. The first group consists of models that are fine-tuned based on the tool dataset, represented by ToolLLaMA Qin et al. (2024). The second group encompasses untrained general language models such as Vicuna Chiang et al. (2023), and Mistral-Instruct-7B Jiang et al. (2023). The third category represents the closed-source model, embodied by GPT4. We subsequently contrast our approach with two agent frameworks, ReACT Yao et al. (2022) and DFSDT Qin et al. (2024), both of which are utilized for multi-step reasoning and model invocation. Notably, all methods employ the ground truth toolset for tool selection, thereby eliminating the influence of the tool retriever.
4.3 Main Experiments
@@
Table 3 displays the results of the comprehensive evaluation of our proposed framework on ToolBench. For the untrained LLMs, it is clear that existing agent frameworks do not improve their performance in tool planning tasks; Vicuna, and Mistral-Instruct-7B all failed at the given tasks with the ReACT and DFSDT frameworks. However, Smurfs exhibits exceptional performance: Mistral combined with Smurfs achieves the highest score among the baselines. Through its task decomposition mechanism, Smurfs transforms a long-context task into several simpler tasks, enabling the untrained model to effectively utilize external tools to manage complex tasks. Regarding the closed-source models, specifically GPT4 in these experiments, Smurfs also demonstrates competitive performance on the benchmark compared to other agent frameworks. A high win rate suggests that Smurfs is more adept at finding better solution paths than ChatGPT, a success that can likely be credited to the verifier agent.
4.4 Ablation Study
4.4.1 Importance of each component in MAS
Average
Pass Win
Mistral 0.0 0.0
Mistral with Smurfs 79.3 86.6
   w/o Verifier Agent $74.5_{\downarrow 6.0\%}$ $83.8_{\downarrow 3.2\%}$
   w/o Answer Agent $73.3_{\downarrow 7.6\%}$ $81.7_{\downarrow 5.7\%}$
   w/o Planning Agent $64.0_{\downarrow 19.3\%}$ $82.9_{\downarrow 4.3\%}$
Table 4: Ablation study on ToolBench I2-Cat., I3-Inst. with Mistral-Instruct-7B.
We performed an ablation study to investigate the impact of each agent in our framework. We removed each agent individually, except for the indispensable Executor Agent, and compared the results to the complete framework. Table 4 shows that the Planning Agent is the most crucial component, followed by the Answer Agent and the Verifier Agent. (1) Verifier Agent Removal: Without verification, the framework uses a general depth-first search, leading to increased computational demand and more tool invocations. (2) Answer Agent Removal: Removing this agent means the Executor Agent’s answer won’t be summarized, risking the ’lost-in-the-middle’ problem due to lengthy tool responses. (3) Planning Agent Removal: Without this agent, the global path searching strategy is affected. Models with Smurfs may show reduced performance without preliminary planning, as seen in current frameworks like ReACT and DFSDT.
4.4.2 Effect of number of agents
Models # of I2-Cat. I3-Inst. Average
Agents Pass Win Pass Win Pass Win
Vicuna 1 49.5 53.0 60.0 86.0 54.8 69.5
2 71.0 66.0 75.0 83.0 73.0 74.5
3 70.5 64.3 78.0 87.0 74.3 75.6
Mistral 1 65.5 70.5 56.0 94.0 60.8 82.3
2 76.0 75.5 73.0 92.0 74.5 83.8
3 79.5 79.2 79.0 94.0 79.3 86.6
GPT4 1 60.5 71.8 57.0 89.0 58.8 80.4
2 69.5 77.8 59.0 94.0 64.3 85.9
3 72.0 77.0 64.0 89.5 68.0 83.3
Table 5: The impact of MAS complexity on the performance of different models. More agents don’t always bring more benefits.
@@ 4.4.2 Effect of number of agents
In addition to evaluating the importance of each component, another relevant question arises: Can a fixed agent system effectively accommodate every model? Our hypothesis suggests that larger language models possess comprehensive intelligence, and employing complex agent systems may potentially hinder their performance compared to simpler ones. As shown in Table 5, we conducted an ablation study on Vicuna, Mistral-Instruct-7B, and GPT4, varying the number of agents ***Agents are incrementally added to the system based on their importance.. The findings indicate that as the number of agents in a MAS increases, leading to increased complexity of the MAS, the system’s performance enhancement does not scale linearly. In certain instances, GPT4 with a 2-agent MAS outperforms the one with a 3-agent MAS, suggesting that larger language models do not consistently benefit from more complex Multi-Agent Systems. Smaller language models can benefit more from the agent system, as seen with Vicuna increasing from 69.5 to 75.6.
@@ 4.5 Case Study
As shown in Figure 4, even though GPT4-DFSDT and GPT4-Smurfs use the same tool calls to solve the problem, GPT4-DFSDT only answers the first sub-question correctly while GPT4-Smurfs answers both sub-questions accurately. In the process of addressing the second sub-question, it is notable that the tool response only mentions titles of film and television products related to "Star Wars", without addressing OTT platforms. GPT-4-DFSDT erroneously interprets these titles as responses to the question, while GPT-4-Smurfs adeptly identifies this discrepancy and provides a more appropriate response. This case highlights that in situations where tool responses are lengthy and questions are complex, the single agent framework like DFSDT may be susceptible to distractions from irrelevant information, leading to erroneous answers. Conversely, the context-efficient Smurfs framework demonstrates a reduced susceptibility to irrelevant information, thereby generating more accurate answers.
5 Related Work
In this section, we review the literature related to multi-agent collaboration, tool-augmented language models, and complex task planning and reasoning. Each subsection provides an overview of the topic, discusses recent advancements, identifies challenges, and suggests avenues for future research.
5.1 Multi-Agents Collaboration
Multi-agent systems have garnered significant attention due to their applicability in various domains such as robotics, economics, and computer networking. One of the key challenges in multi-agent systems is achieving effective collaboration among autonomous agents. Recent studies have focused on understanding collaborative behaviors and developing coordination strategies to enable agents to work together towards common goals. For example, research in reinforcement learning has explored techniques for emergent coordination, where agents learn to collaborate through interaction with the environment and other agents. Du et al. (2023); Liang et al. (2023) carry out multi-agent interaction in the form of debate, which improves mathematical and strategic reasoning tasks. Li et al. (2023); Wang et al. (2023) uses role-playing to conduct the interaction between multi-agents.
5.2 Tool-Augmented Language Models
Language models augmented with external knowledge sources, often termed as tool-augmented language models, have demonstrated promising outcomes in a variety of natural language processing tasks. These models utilize external knowledge bases, ontologies, or pre-trained models to enhance their contextual understanding and boost performance in tasks such as text generation, summarization, and question answering. Recent advancements in this field include techniques for integrating structured knowledge into language models, such as graph-based representations or semantic parsing. For instance, Qin et al. (2024) introduces ToolBench, an instruction-tuning dataset for tool usage, along with a fine-tuned tool-oriented model, ToolLLaMA. Another significant contribution is Gorilla Patil et al. (2023), which excels at writing API calls and also introduces a benchmark for evaluating LLMs with tools.
5.3 Complex task planning and reasoning
Problem decomposition (Zhou et al. (2023); Drozdov et al. (2022); Khot et al. (2022); Press et al. (2022)) is a popular paradigm used in the LLM and leads to good performance in challenging reasoning tasks. This divide-and-conquer strategy provides sufficient explanation of how the model works. The current trend is that when model size reaches some limit, people start looking for a way to make the most of the limited model. Chain-of-thought prompting Wei et al. (2022) is one of the strategy that lead models to divide the tasks into a chain, and solve the task step by steps. Chen et al. (2022) propose a program of thoughts structure to enhance the reasoning ability. Yao et al. (2024) simulate thought as tree structure to solve task. Besta et al. (2024) using graph to simulation.
@@ 6 Conclusion
In this study, we present a novel MAS framework, Smurfs, tailored to enhance the planning and reasoning capabilities of LLMs in handling complex tasks that involve lengthy contexts and tools. We conducted experiments on the multi-step tool usage benchmark, ToolBench, and the results demonstrated the overall effectiveness of the Smurfs framework compared to the baseline models.
Ablation studies are carried out to investigate and compare the significance of different components within the MAS framework. The findings revealed that preliminary planning was the most crucial element. Content summarizing also played a key role in mitigating the ’lost-in-the-middle’ issue often encountered in long-context multi-step reasoning scenarios. While verification was not as influential on effectiveness, it proved valuable in enhancing computational efficiency by identifying the optimal solution path for complex task resolution. By dissecting and comparing the different aspects of the MAS framework, we aim to offer insights that could inspire advancements in the applicability and accuracy of LLMs.
In conclusion, this research contributes to the expanding field of study focused on enhancing LLM capabilities, particularly for multi-step tool usage tasks. It emphasizes the importance of task decomposition, preliminary planning, and efficient verification for improving task execution performance. We are confident that the knowledge gained from this study will lay the groundwork for the development of more sophisticated and efficient LLM frameworks in the future.

### v3

@@ 3.4 Subtask Solving Process
The Executor Agent is responsible for choosing and executing the tools to solve the sub-tasks. At each time t, the agent can invoke one tool to tackle the given task:
$\gamma=EA.gen\_thought(p,M,\tau,h)$ (2)
$\alpha=EA.choose\_tool(p,\gamma,\tau)$ (3)
$\beta=EA.gen\_arguments(p,M,D[\alpha])$ (4)
$r=EA.call\_tool(\alpha,\beta)$ (5)
Where p is the sub-problem from Planning Agent, h is the hint from the Verifier Agent, $\tau$ is the tool list, M is local memory, $D[\alpha]$ means the tool document of tool $\alpha$. The agent, using the ReACT format Yao et al. (2022) to choose the tool and arguments, then execute the tool. Noticed that each inference process only uses the relevant part from the local memory and tool list to reduce the context redundancy. More detailed information of the Executor Agent can be found in Figure 6.
@@
To mitigate the performance degradation caused by lengthy contexts, we introduce the Answer Agent role, designed to extract crucial content for each step and sub-problem:
$Answer:a=AA(q,r,M)$ (6)
Where q is sub-problem from the Planning Agent, r is response from the Executor Agent, M is the local memory (or global memory if max retry reaches). As the ‘lost-in-the-middle’ theory described in section 2.1, retaining all information may not always be beneficial, particularly in cases where the solution path is challenging to discern. Therefore, the primary role of the Answer Agent is to succinctly summarize the generated answers and tool responses to maintain the memory efficiency.
@@
The Verifier Agent serves as an early-stopping and reflection mechanism, allowing for a balance between effectiveness and efficiency:
$h,c=VA(q,a)$ (7)
Where q denotes the sub-problems from the Planning Agent, a denotes the answer from the answer agent, h and c denotes hint and check status respectively. If check status generated is 0, that means the Verifier Agent thinks the sub-problem isn’t completed, the system will add the thought and answer of this time to the local and global memory, set t=t+1 and continue the inference procedure.If check status is 1, the sub-problems is thought to be solved and the system will start to deal with the next sub-problem.
@@ 4 Experiments
To evaluate both the effectiveness and efficiency of the Smurfs framework, in thie section, we carried out two multi-tool planning tasks: (1) an open-ended task, StableToolBench Guo et al. (2024), and (2) a closed-ended task, HotpotQA Yang et al. (2018). In addition to these main experiments designed to assess the entire framework, we conducted an ablation studies followed by a case study to test the capabilities of each component within the multi-agent framework and investigate the underlying reasons for its effectiveness.
4.1 Open-ended Task: StableToolBench
@@
StableToolBench is a tool learning benchmark derived from ToolBench Qin et al. (2024), encompassing multi-step tool usage tasks across over 16,000 APIs. The benchmark employs two metrics for evaluation: (1) Pass Rate measures the percentage of instructions successfully executed within the allocated budget. (2) Win Rate represents the preference selection by a ChatGPT evaluator when presented with two solution paths.
Baselines
Following the original paper that introduced the benchmark, we adopt ReACT (CoTWei et al. (2023) and DFSDT Touvron et al. (2023) as baseline methods for comparison. Additionally, we include the backbones used in the paper: gpt-3.5-turbo-0613 (GPT-3.5 Turbo) OpenAI and gpt-4-turbo-preview (GPT-4 Turbo). To explore the adaptability of the tool-planning methods, we also include Mistral-7B-Instruct-v0.2 (Mistral-7B) Jiang et al. (2023) as one of the selected backbones in our experiments.
Settings
To minimize the influence of varying tool APIs on experimental results, we conducted all experiments using the same API cache Guo et al. (2024). For a fair comparison among the candidate methods and to reduce variability, each model was executed once and evaluated three times, with results averaged. Other settings follow those specified in the original benchmark paper.
Results
@@
Table 3 displays the results on StableToolBench. For the untrained LLM, Mistral-7B, existing agent frameworks did not improve its performance in tool planning tasks; Mistral-7B failed these tasks when integrated with the ReACT and DFSDT frameworks ***Experiment results show that Mistral-7B failed to correctly execute the ‘finish’ action during inference, resulting in invalid responses.. However, Smurfs exhibited exceptional performance: when combined with Mistral-7B, Smurfs achieved competitive scores among the baselines. Through its task decomposition mechanism, Smurfs transforms long-context tasks into simpler ones, enabling the untrained model to effectively utilize external tools for managing complex tasks. Regarding closed-source models, specifically GPT4 in these experiments, Smurfs also demonstrated outstanding performance on the benchmark compared to other agent frameworks and achieved state-of-the-art results on the benchmark. Its high success rate suggests that Smurfs is more effective at finding optimal solution paths compared to ChatGPT.
Further Analysis
We conducted a detailed analysis of the token costs associated with each tool planning method for the tasks, a critical evaluation aspect for multihop reasoning tasks. As shown in Table 1 (detailed in Appendix E), the average token costs per question and API request are evaluated for ReACT, DFSDT, and Smurfs on StableToolBench. The analysis reveals that DFSDT generally requires about 20,000 tokens per question, encompassing both prompt and completion tokens. This is nearly three times the token cost compared to ReACT and twice as much as Smurfs. Despite this higher token cost, DFSDT does not demonstrate commensurate effectiveness improvements over other methods. These findings underscore the cost-efficiency of the proposed MAS framework, Smurfs, which not only reduces token expenditure in solving multihop planning tasks but also delivers outstanding performance in evaluations.
[…]
Easy Medium Hard All
[…]
4.2 Closed-ended Task: HotpotQA
Compared to open-ended tasks, closed-ended tasks provide a more stable and robust evaluation. To this end, we evaluate the methods on HotpotQA Yang et al. (2018) in addition to StableToolBench. HotpotQA is a multi-hop QA task that is challenging due to the requirement for rich background knowledge, with answers typically being short entities or yes/no responses.
Baselines
The compared baselines include CoT Wei et al. (2023), ReActYao et al. (2022), ChameleonLu et al. (2023), Reflexion Shinn et al. (2023), BOLAA Liu et al. (2023), ReWOO Xu et al. (2023), FireAct Chen et al. (2023a), AutoActQiao et al. (2024).
Settings and Metrics
Following the settings in  Qiao et al. (2024), we use open-source Llama-2 models Touvron et al. (2023) and Mistral-7B Jiang et al. (2023) as the backbones of each agent to evaluate the performance of Smurfs. The evaluation metrics is $\texttt{reward}\in[0,1]$, defined as the F1 score grading between the prediction and ground-truth answer. For more details about the experiment, see Appendix C.
Results
Smurfs, as an untrained MAS system, not only comprehensively outperforms untrained agents but also achieves and even surpasses the accuracy of trained agents across most backbone models. This sufficiently demonstrates that the mechanism of smurfs ensures strong generalization capabilities while maintaining high effectiveness.
Observations indicate that the performance of LLama-2-13b-chat on smurfs-related tasks is suboptimal, likely due to its limited capabilities in tool arguments generation. Specifically, the primary issue identified is that, when the Executor agent successfully selects relevant tool, it tends to produce hallucination arguments that can’t be used by the tools. This indicates that LLama-2-13b-chat may need further training for usage of tools. The experimental results may substantiate this viewpoint, demonstrating that the untrained methods of llama-2-13b-chat generally exhibit significantly lower accuracy compared to the trained methods. Nevertheless, Smurfs achieves the second highest accuracy among the untrained methods, only slightly behind reflexion, which still attests to Smurfs’ capability.
I3-Inst.
[…]
GPT-3.5 Turbo with Smurfs $60.1_{\pm 1.0}$ 57.4
   w/o Answer Agent $57.4_{\pm 2.9}$ $49.2$
   w/o Verifier Agent $54.1_{\pm 2.7}$ $42.6$
   w/o Planning Agent $35.5_{\pm 3.3}$ $42.6$
GPT-4 Turbo with Smurfs $70.5_{\pm 1.0}$ 72.1
   w/o Answer Agent $82.2_{\pm 2.5}$ $72.1$
   w/o Verifier Agent $79.2_{\pm 0.8}$ $63.9$
   w/o Planning Agent $71.9_{\pm 2.8}$ $63.9$
Table 5: Ablation study on StableToolBench I3-Inst subset to investigate the importance of each component within the framework.
@@
5 Ablation Study
5.1 Importance of each component in MAS
We performed an ablation study to investigate the impact of each agent in our framework. We removed each agent individually, except for the indispensable Executor Agent, and compared the results to the complete framework. Table 5 shows that the Planning Agent is the most crucial component, followed by the Verifier Agent, with the Answer Agent being the least important.
(1) Verifier Agent Removal: Without verification, the framework uses a general depth-first search, leading to increased computational demand and more tool invocations.
(2) Answer Agent Removal: Removing this agent means the Executor Agent’s answers won’t be summarized, risking the ’lost-in-the-middle’ problem due to lengthy tool responses. As shown in the results, a more intelligent model, GPT-4 Turbo, can mitigate the negative impact of the Answer Agent’s removal. We believe this is because the more powerful model can leverage more information effectively.
(3) Planning Agent Removal: Removing the Planning Agent affects the global path-searching strategy. Models with Smurfs may show reduced performance without preliminary planning, as seen in current frameworks like ReACT and DFSDT. The results demonstrate that the impact of removing the Planning Agent is significant, as it directly influences the multihop reasoning ability of the MAS.
@@ 5.2 Case Study
As shown in Figure 3, although GPT4-DFSDT and GPT4-Smurfs use the same tool calls to solve the problem, GPT4-DFSDT only answers the first sub-question correctly while GPT4-Smurfs answers both sub-questions accurately. In the process of addressing the second sub-question, it is notable that the tool response only mentions titles of film and television products related to "Star Wars", without addressing OTT platforms. GPT-4-DFSDT erroneously interprets these titles as responses to the question, while GPT-4-Smurfs adeptly identifies this discrepancy and provides a more appropriate response. This case highlights that in situations where tool responses are lengthy and questions are complex, the single agent framework like DFSDT may be susceptible to distractions from irrelevant information, leading to erroneous answers. Conversely, the context-efficient Smurfs framework demonstrates a reduced susceptibility to irrelevant information, thereby generating more accurate answers.
@@ 6 Conclusion
In this study, we present a novel MAS framework, ‘Smurfs’, tailored to enhance the planning and reasoning capabilities of LLMs in handling complex tasks that involve lengthy contexts and tools. We conduct experiments on the multi-step tool usage benchmark, StableToolBench and HotpotQA, and the results demonstrate the overall effectiveness and efficiency of the Smurfs framework compared to baseline methods.
In conclusion, this research contributes to the expanding field of study focused on enhancing LLM capabilities, particularly for multi-step tool usage tasks. It emphasizes the importance of task decomposition, preliminary planning, and efficient verification for improving task execution performance. For future work, we believe incorporating more dedicated and specific roles within the system may further enhance effectiveness and efficiency, based on the ‘Smurfs principle’: synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs.